# MicroEvoEval: A Systematic Evaluation Framework for Image-Based Microstructure Evolution Prediction

### v1

@@ Appendix D Implementation Details of Our Benchmark
Here, we will further introduce all the implementation details. The corresponding data and code can be found in the supplementary materials. Due to the extensive size of our processed dataset (79G), we are only able to upload a sampled test dataset for each task in the supplementary materials. We promise that our full datasets and the code of our benchmark will be made publicly available on Hugging Face and GitHub after official publication with cc-by-nc-4.0 and Apache-2.0 license, respectively. We are committed to maintaining and updating them based on community feedback and our future research.
Please note: To ensure the anonymity of the peer review process, no information identifying the authors of this paper has been included in the appendix.
@@ D.1 Implementation Details of Datatsets
Here, we show some data instances from the training (10-10) and test (10-90) sets of the four datasets in our benchmark (Figure 8 - 15). Some intermediate frames are omitted for brevity.
@@
PredRNNv2 (wang2022predrnn) TPAMI 2022
@@ D.2 Implementation Details of Metrics
In our benchmark, we totally have five numerical accuracy metrics, two physical fidelity metrics and the computational efficiency in the main text and Appendix.

### v2

@@ Appendix D Implementation Details of Our Benchmark
Here, we further introduce all the implementation details. The corresponding data and code can be found in the main paper. Our full datasets and the code of our benchmark are publicly available on Hugging Face and GitHub with cc-by-nc-4.0 and Apache-2.0 license, respectively. We are committed to maintaining and updating them based on community feedback and our future research.
@@ D.1 Implementation Details of Datatsets
Here, we show some data instances from the training (10-10) and test (10-90) sets of the four datasets in our benchmark (Figure 12 - 19). Some intermediate frames are omitted for brevity.
@@
PredRNN.v2 (wang2022predrnn) TPAMI 2022
@@ D.2 Implementation Details of Metrics
In our benchmark, we totally have four numerical accuracy metrics, two physical fidelity metrics and the computational efficiency in the main text and Appendix.
@@
D.3 Implementation Details of Dumbbell Dataset
In this part, we describe the raw data and the metrics we used in this new Dumbbell dataset. Unless otherwise specified, other details will remain consistent with the other four datasets.
The raw data.
The statistic of the train dataset, validation dataset and test dataset are (200, 100, 64, 64), (30, 100, 64, 64) and (50, 100, 64, 64) respectively. We split the train dataset and the validation dataset into $1800$ samples (each containing 20 frames) and 270 samples (each containing 20 frames) respectively. The test dataset is splitting into 450 samples (each containing 20 frames) for short-term prediction (10-10), 200 samples (each containing 60 frames) for long-term prediction (10-50) and 50 samples (each containing 100 frames) for long-term prediction (10-90).
OOD dataset construction.
To evaluate the models’ out-of-distribution generalization ability, we additionally construct an OOD version of the Dumbbell dataset with the same size of standard version by shifting the physical parameters radius. In the meta-dataset used for training and validation, the dumbbell radius is sampled from the interval $[0.31,0.39]$, in contrast, the OOD dataset is generated using a larger radius range of $[0.39,0.43]$, which lies entirely outside the training distribution.
The metrics.
In this dataset, we also have the four numerical accuracy metrics that we describe in Appendix D.2. Due to the unique physical properties of this new dataset, we also plan to design new, corresponding physical metrics in our future work.